IF the slopes of the lines 2y-6bx = 7 & 5y+30x+13 = 0 are equal, find the value of b?.
Answer by grishma.kshatriya(27)

Solution :
Consider the first equation as 2y-6bx = 7
Write this equation in the slope intercept form
y = mx + b where m is the slope ,
2y - 6bx = 7
Adding 6bx on both the sides ,
2y = 6bx + 7
Divide by 2 on both the sides ,
y = 3bx + 7/2
So here comparing with the y=mx + b ,
m1 = 3b
Similarly the second equation 5y+30x+13 = 0
subtract the 30x and 13 from both the sides we get ,
5y = -30x - 13
Divide both the sides by 5
y = -6x - 13/5
So by comparing m2 = -6
Now it says that the slopes of the lines are equal
m1 = m2
3b = -6
b = -2
Thank you .